OpenStudy (kaylala):

proving identities: tan - sec = -(cos/(1+sin))

OpenStudy (usukidoll):

@UnkleRhaukus

OpenStudy (loser66):

(sin/cos)-(1/cos)= (sin-1)/cos time and divided by (sin +1) (sin^2 -1)/(cos ( sin +1) = (-cos^2)/(cos (sin+1) =(-cos)/(1+sin)

OpenStudy (kaylala):

thanks @Loser66